#1: What is the thinnest soap film that will strongly reflect red light from a helium-neon laser? The wavelength of this light is 633 nm in air and 470 nm in soapy water.

#2 What is the wavelength of the light from a carbon dioxide laser when it is in glass with an index of refraction of 1.7? The wavelength in a vacuum is 10600 nm.
Answer #1

1)What is the thinnest soap film that willstrongly reflect red light from a helium-neon laser? The wavelengthof this light is 633 nm in air and 470 nm in soapy water.(nm)

{Min Thickness} = (1/4)*(?_film) = (1/4)*(470 nm) = 117.5 nm